LUCY MAKES A RECTANGLE WITH A PERIMTER OF 18CM THE LENGTH IS TWICE THE WIDTH CALACULATE THE AREA OF THE RECTANGLE
Found 2 solutions by vasumathi, amoresroy:
Answer by vasumathi(46)

The perimetre = P = 18 cms (given)
Let the width be W so the length would be 2W
perimetre in terms of L and W is ...
P = 2(2W + W)
P = 2*3W
18 = 6W
now divide both sides by 6
W = 18/6 = 3
L = length = 2 W = 2*3 =6
Area of the rectangle = Length * width = 6* 3 = 18 square cms
This is the final answer

Answer by amoresroy(361)

LUCY MAKES A RECTANGLE WITH A PERIMTER OF 18CM THE LENGTH IS TWICE THE WIDTH CALACULATE THE AREA OF THE RECTANGLE
Let w = the width of the rectangle
2w = the length of the rectangle
2(w+2w) = 18
2(3w) = 18
w = 3
Area = w*2w
= 3*2*3
= 18
The area of the rectangle is 18 sq. cm.